# The Impact of Domain Knowledge and Multi-Modality on Intelligent Molecular Property Prediction: A Systematic Survey

### v1

@@ 7 Conclusion
In this paper, we discuss the significant role of multi-modal data and domain knowledge in enhancing molecular property prediction through DL methods. We explored various molecular data formats and domain knowledge, crucial in understanding molecular complexities. Our review of recent encoder architectures and training strategies highlighted how integrating domain knowledge and multi-modal data advances these models. By benchmarking prominent works, we provided a comparative analysis of their effectiveness. Ultimately, our discussion revealed the profound impact of domain knowledge and multi-modal data in DL approaches, marking a transformative advancement in drug discovery and computational molecular analysis.

### v3

@@ 7 Conclusion
In this paper, we discuss the significant role of multi-modal data and domain knowledge in enhancing molecular property prediction through DL methods. We explored various molecular modalities and domain knowledge, crucial in understanding molecular complexities. Our review of recent encoder architectures and training strategies highlighted how integrating domain knowledge and multi-modal data advances these models. By benchmarking prominent works, we provided a comparative analysis of their effectiveness. Ultimately, our discussion revealed the profound impact of domain knowledge and multi-modal data in DL approaches, marking a transformative advancement in drug discovery and computational molecular analysis.